In a standing wave pattern in a vibrating air column, nodes are formed at a distance of 4.0 cm. if the speed of sound in air is 328 m/s, what is the frequency of the source?

Grade:12

1 Answers

Arun
6 years ago
Here given that 1 = 50 cm, v = 340 m/s As it is an open organ pipe, the fundamental frequency f base 1 = (v/21) = 340/2 * 50 * 10^-2 = 340 Hz. So, the harmonies are f base 3 = 3 × 340 = 1020 Hz f base 5 = 5 × 340 = 1700, f base 6 = 6 × 340 = 2040 Hz So, the possible frequencies are between 1000 Hz and 2000 Hz are 1020, 1360, 1700